# Finding Flawed Fictions: Evaluating Complex Reasoning in Language Models via Plot Hole Detection

### v2

@@ 8 Conclusion
While FlawedFictionsMaker offers a general approach for generating continuity errors, future work could explore methods providing finer control over the types and complexity of introduced plot holes. Additional research might focus on designing new post-training strategies that can enhance model performance on FlawedFictions. Another promising direction would be to investigate whether using FlawedFictionsMaker to generate large amounts of synthetic training data could enhance LLMs’ reasoning capabilities more broadly. Future work can also consider plot deficiencies other than plot holes, like plot conveniences or coincidences (termed cheap plot tricks Ryan (2009)) or apply similar approaches to non-fictional contexts like fact-checking, misinformation detection, and education.
@@ A.1 A More Formal Treatment of Continuity Errors
We discussed in §2 that the Definition 2.1 fails to account for implicit knowledge such as our world understanding and beliefs that are often essential to reason about contradictions in stories. We utilize the Possible Worlds theory from Lewis (1978) to extend our definition. The core contribution of Lewis’s theory is to assess truthfulness of the statements that are never stated in the text of the narrative. E.g. can we say that Sherlock lived closer to Paddington Station than Waterloo Station? While using a map of real world London one can check Baker Street being closer to Paddington Station, story’s text never explicitly states this. However, we can still assign truth to this statement since we do not have any special reason to believe that geography of London in Sherlock Holmes is remarkably different from the real world. To decide if a proposition $p$, which is true in the belief world of the reader (or community of readers) is also true in story $f$$\text{isTrue}(f,p)$—, without explicitly being stated in $f$, Lewis (1978) uses the notion of counterfactuals. Specifically, a proposition $p$ is non-vacuously true in $f$, when some world where $f$ is told as fact and $p$ is true, is closer to the belief world of the reader $W_{b}$, than any world where $f$ is told as fact and $p$ is not true. Hence, while we can consider a world where Sherlock Holmes is told as fact and London is arranged very different from the real world such that Baker Street is closer to the Waterloo Station than Paddington Station, that world will be further away from the belief world of the reader compared to a world that preserves the geography of London.
@@
In other words, if using all the propositions in $\mathcal{F}$ except $\phi_{e}$, as well as the propositions from the belief world that are non-vacuously true in $f$ [^8] Here, $f$ is a story $f^{\prime}$ where $\phi_{e}$ is never stated., we can conclude that the negation of $\phi_{e}$ is true, that means $\phi_{e}$ represents a continuity error in $f$.
According to the possible worlds theory, stories $f$ with such logical contradictions lead to impossible fictions, where there exists no possible world where the story is told as fact, i.e. $\mathcal{W}_{f}=\{\}$. In principle, for such impossible story, any statement $p$ is vacuously true. However, such a treatment can be too harsh especially when the logical contradictions are accidental and not blatantly renders the plot useless (e.g. we can still make sense of a story even if a wound placement on a character has changed without notice). There are formalizations to non-vacuously evaluate truth statements in impossible worlds in Lewis (1978) and follow-up work Alber (2019); Badura & Berto (2019), however that falls out of the scope of this work. Our primary concern here is understanding if LLMs can reason when a story represents worlds that are impossible.
@@
Continuity Error Sherlock Holmes by Sir Arthur Conan Doyle When we are first introduced to Watson in A study in pink, he is described as having injury in his left arm, but the very next story A sign of Four contradicts this where his war wound is on his knee. Harmless Lewis (1978)
Citizen Kane (1941) In the film Kane dies alone, but a group of reporters are trying to discover meaning of his dyning words. If he died alone who heard the words Rosebud? Harmless Ryan (2009) Example of incorporating real world beliefs to reason about plot holes – “when people die alone that means no one could hear their last words” is a proposition we know to be true from our commonsense and not something stated in the story
Out of Character Behavior Little Red Riding Hood by Brothers Grimm A mother tells her daughter, Little Red Riding Hood, to go through the forest and to bring some food to her ailing grandmother. She warns the little girl not to talk to strangers. On her way, Little Red Riding Hood meets a hungry wolf and tells him about her mission. The wolf runs to the grandmother’s house, eats her, and takes her place in bed. When Little Red Riding Hood arrives she mistakes the wolf for the grandmother. After a conversation during which he pretends to be the grandmother, the wolf jumps out of the bed and eats Little Red Riding Hood. Why did he not just eat her when they met for the first time? Unbridgeable Ryan (2009)
@@
Impossible Event Dark Knight Rises (2012) In The Dark Knight Rises (2012), a full team of police members was trapped underground for months, yet they all walk out cleanshaven and well-dressed. Harmless Davids (2022)
@@ A.2 Human Annotation and Benchmarking
The annotators were hired from the Prolific [^9] https://app.prolific.com/ platform with the screening conditions that the candidates have English as their primary language, are residents of UK, US, or Canada, have at least an undergraduate degree, and face no literary difficulties. We also conducted a screening test where candidates were given a small set of examples from the task for which the ground truths were already verified by the authors and selected candidates for the actual study who performed well on this screening test. The selected examples had 50% samples that were incorrectly assessed by ChatGPT and we made use of this to find candidates who were potententially using LLMs for annotations. We also checked the average amount of time it took for participants to complete the pilot study, and didn’t consider those who solved the task too quickly, with the risk of them potentially using LLMs. We finally ended up recruiting 19 annotators, who were paid \$12 per hour for their work with extra 20-30% bonuses each time they annotated more than 10 stories. Estimated time per annotation for each example was 5 minutes and we ended up paying a total of \$$\$6500$ to the annotators. We got roughly 350 stories annotated, and got at least 3 annotations for each story. An example of our annotation framework built using Argilla [^10] https://github.com/argilla-io/argilla is provided in Figure 3.
@@
We recruited 9 undergraduates with English major and present them with the same task of plot hole detection and the same specifications and instructions as we do for different LLMs. We sampled 50 examples from our dataset and obtained 3 responses for each instance. The estimated time for solving each task was 15 minutes (approximated by the first author) and participants were compensated \$5 for providing response for each story, thereby providing \$20 per hour for their work. To encourage participants to give their best efforts towards solving the task, we provide a 30% bonus for solving the task with higher accuracy ($>70\%$ accuracy on the classification task). We paid a total of \$944.60 to the participants. An example of the interface has been provided in Figure 4. The complete study document shared with the participants is included at the end of this paper §A.11.
@@ A.4 More Details on Experimental Setup
We augment the plot hole detection model i.e. generator with a verifier model (Cobbe et al., 2021) that validates if the plot hole detected by the generator is legitimate. If it is deemed illegitimate, we sample from the generator again, till either the verifier agrees or generator answers by saying No continuity error detected. The maximum number of samples from the generator are capped at 5. For the verifier we use Claude 3.5 Sonnet model prompted to test the validity of a proposed plot hole. Due to increased cost with using a verifier we only report results when Claude 3.5 Sonnet generator is augmented with the verifier.
@@ A.5.2 Factors Effecting Performance on FlawedFictions
We investigate if length of a story has an effect on how accurately do different LLMs detect continuity errors in them by measuring correlation  [^11] We use Point-Biserial Correlation since CEEval-Full at an instance level is a discrete i.e. 0 or 1. between a story’s length (measured by counting number of words) and the CEEval-Full score on that story. We find negative correlation coefficients for all the models that we test and while the correlation values are low -0.1 to -0.2, for 13 out of 14 models the correlation observed is statistically significant (p-value $<$ 0.05). Refer to the Table 7 for the exact values.
@@ A.9.1 Summarization
Summary generated with GPT-4o-mini Once upon a time, there was a poor woman living in a wretched hut, whose little boy ventured into the cold forest to gather wood. While wandering, he discovered a jagged white stone and, feeling pity for it, draped his jacket over it. Upon returning home without his jacket, his mother scolded him for his foolishness, insisting that charity begins at home. When the boy returned to the stone, he found it had turned over, revealing a chest filled with silver coins. Suspecting the money to be stolen, he tossed it into a nearby pond, keeping only a four-shilling piece that floated, believing it to be honest. The boy’s mother, furious at his actions, sent him away to seek his fortune. He wandered far and wide, but was deemed too small and weak for any work until he found a merchant who hired him to fetch wood and water. When the merchant prepared to journey to distant lands, the boy entrusted him with the four-shilling piece to buy something in return. The merchant, initially reluctant, ended up purchasing a cat from an old woman who wished to dispose of it. As the merchant sailed, a fierce storm struck, and upon reaching a strange land, he found a tavern filled with mice, where diners struggled to eat. Remembering the cat, he brought it forth, and the patrons were able to enjoy their meal in peace. They begged him to sell the cat, and he agreed for a hundred dollars. Afterward, the merchant encountered another storm and, upon reaching yet another land, sold the cat again for two hundred dollars, and then once more for three hundred dollars in a place plagued by larger rats. Throughout his journey, the merchant pondered how much of his earnings to give the boy, deciding initially to keep a portion for himself. However, when a terrible storm threatened the ship, he vowed to give the boy all the money. Upon reaching home safely, he honored his promise, presenting the boy with six hundred dollars and his daughter as well. The boy, once a mere scullion, became as wealthy as the merchant, living in splendor and happiness, and graciously took in his mother, declaring that charity does not solely begin at home.   Claude 3.5 Sonnet with Verfier’s Explanation of the Continuity Error There is a continuity error in the story regarding the amount of money the merchant earned from selling the cat. The merchant sold the cat three times for progressive amounts: $100,$200, and $300,whichwouldtotal$600. However, this creates a continuity error because once the cat was sold the first time, the merchant would no longer have possessed it to sell it two more times. Unless the story explicitly explained how the merchant regained possession of the cat between sales, this represents a logical impossibility in the narrative sequence.   Author’s Remark. The original story has the cat return to the ship every time a sale is done, a detail which is missing in the summary that creates the continuity error if cat was sold once, how the merchant managed to sell it again and again.

### v3

@@ 8 Conclusion
While FlawedFictionsMaker offers a general approach for generating continuity errors, future work could explore methods providing finer control over the types and complexity of introduced plot holes. Additional research might focus on designing new post-training strategies that can enhance model performance on FlawedFictions. Another promising direction would be to investigate whether using FlawedFictionsMaker to generate large amounts of synthetic training data could enhance LLMs’ reasoning capabilities more broadly. Future work can also consider plot deficiencies other than plot holes, like plot conveniences or coincidences (termed cheap plot tricks ryan2009cheap) or apply similar approaches to non-fictional contexts like fact-checking, misinformation detection, and education.
@@ A.1 A More Formal Treatment of Continuity Errors
[…]
@@
In other words, if using all the propositions in $\mathcal{F}$ except $\phi_{e}$, as well as the propositions from the belief world that are non-vacuously true in ff [^8] Here, $f$ is a story $f^{\prime}$ where $\phi_{e}$ is never stated., we can conclude that the negation of $\phi_{e}$ is true, that means $\phi_{e}$ represents a continuity error in $f$.
According to the possible worlds theory, stories $f$ with such logical contradictions lead to impossible fictions, where there exists no possible world where the story is told as fact, i.e. $\mathcal{W}_{f}=\{\}$. In principle, for such impossible story, any statement $p$ is vacuously true. However, such a treatment can be too harsh especially when the logical contradictions are accidental and not blatantly renders the plot useless (e.g. we can still make sense of a story even if a wound placement on a character has changed without notice). There are formalizations to non-vacuously evaluate truth statements in impossible worlds in lewis1978truth and follow-up work alber2019; Badura02012019, however that falls out of the scope of this work. Our primary concern here is understanding if LLMs can reason when a story represents worlds that are impossible.
Below we provide a detailed example to further clarify the definition:
Consider this story:
  • Jack is drinking hot tea in a cup
  • Jack gives his tea to Jill
  • Jill drinks the cold tea from Jack’s cup
From the story we can extract propositions: $\phi_{1}=$ ”Jack drinks tea”, $\phi_{2}=$”tea is hot”, $\phi_{3}=$ ”Jack gives tea to Jill”, $\phi_{4}=$ ”Jill drinks tea”, and $\phi_{5}=$ ”tea is cold”. Note that removing $\phi_{5}$ from our proposition set does not allow us to conclude $\neg\phi_{5}$ (”tea is not cold”) from the remaining propositions alone, since ”hot” does not logically entail ”not cold” without additional assumptions. This is precisely why we need the extended definition with reader beliefs. A reader’s commonsense knowledge includes the belief that ”if tea is hot, then tea is not cold” ($\phi_{2}\to\neg\phi_{5}$). When we include this reader belief in our reasoning, we can derive the contradiction: $\{\phi_{2}\to\neg\phi_{5},\phi_{2}\}\vdash\neg\phi_{5}$ (by modus ponens) This shows that $\phi_{5}$ is logically inconsistent with the rest of the story when combined with reasonable reader beliefs, making it a continuity error according to our extended definition. The key insight is that continuity errors in fiction cannot be detected purely from explicit textual propositions—they require the integration of implicit world knowledge that readers bring to their understanding of the story.
@@
Continuity Error Sherlock Holmes by Sir Arthur Conan Doyle When we are first introduced to Watson in A study in pink, he is described as having injury in his left arm, but the very next story A sign of Four contradicts this where his war wound is on his knee. Harmless lewis1978truth
Citizen Kane (1941) In the film Kane dies alone, but a group of reporters are trying to discover meaning of his dyning words. If he died alone who heard the words Rosebud? Harmless ryan2009cheap Example of incorporating real world beliefs to reason about plot holes – “when people die alone that means no one could hear their last words” is a proposition we know to be true from our commonsense and not something stated in the story
Out of Character Behavior Little Red Riding Hood by Brothers Grimm A mother tells her daughter, Little Red Riding Hood, to go through the forest and to bring some food to her ailing grandmother. She warns the little girl not to talk to strangers. On her way, Little Red Riding Hood meets a hungry wolf and tells him about her mission. The wolf runs to the grandmother’s house, eats her, and takes her place in bed. When Little Red Riding Hood arrives she mistakes the wolf for the grandmother. After a conversation during which he pretends to be the grandmother, the wolf jumps out of the bed and eats Little Red Riding Hood. Why did he not just eat her when they met for the first time? Unbridgeable ryan2009cheap
@@
Impossible Event Dark Knight Rises (2012) In The Dark Knight Rises (2012), a full team of police members was trapped underground for months, yet they all walk out cleanshaven and well-dressed. Harmless essay91967
@@ A.2 Human Annotation and Benchmarking
The annotators were hired from the Prolific [^9] https://app.prolific.com/ platform with the screening conditions that the candidates have English as their primary language, are residents of UK, US, or Canada, have at least an undergraduate degree, and face no literary difficulties. We also conducted a screening test where candidates were given a small set of examples from the task for which the ground truths were already verified by the authors and selected candidates for the actual study who performed well on this screening test. The selected examples had 50% samples that were incorrectly assessed by ChatGPT and we made use of this to find candidates who were potententially using LLMs for annotations. We also checked the average amount of time it took for participants to complete the pilot study, and didn’t consider those who solved the task too quickly, with the risk of them potentially using LLMs. We finally ended up recruiting 19 annotators, who were paid \$12 per hour for their work with extra 20-30% bonuses each time they annotated more than 10 stories. Estimated time per annotation for each example was 5 minutes and we ended up paying a total of $\mathdollar 6500$ to the annotators. We got roughly 350 stories annotated, and got at least 3 annotations for each story. An example of our annotation framework built using Argilla [^10] https://github.com/argilla-io/argilla is provided in Figure 3.
@@
We recruited 9 undergraduates with English major and present them with the same task of plot hole detection and the same specifications and instructions as we do for different LLMs. We sampled 50 examples from our dataset and obtained 3 responses for each instance. The estimated time for solving each task was 15 minutes (approximated by the first author) and participants were compensated \$5 for providing response for each story, thereby providing \$20 per hour for their work. To encourage participants to give their best efforts towards solving the task, we provide a 30% bonus for solving the task with higher accuracy ($70\%$ accuracy on the classification task). We paid a total of \$944.60 to the participants. An example of the interface has been provided in Figure 4. The complete study document shared with the participants is included at the end of this paper §A.11.
@@
For both of the annotation exercises, we got an Institutional Review Board (IRB) exemption from the Human Subjects Division which determined our studies as human subjects research that qualifies for exempt status (Category 101) and the IRB IDs were STUDY00021942 and STUDY00022395 for the verification and benchmarking studies respectively.
@@ A.4 More Details on Experimental Setup
We augment the plot hole detection model i.e. generator with a verifier model (Cobbe2021TrainingVT) that validates if the plot hole detected by the generator is legitimate. If it is deemed illegitimate, we sample from the generator again, till either the verifier agrees or generator answers by saying No continuity error detected. The maximum number of samples from the generator are capped at 5. For the verifier we use Claude 3.5 Sonnet model prompted to test the validity of a proposed plot hole. Due to increased cost with using a verifier we only report results when Claude 3.5 Sonnet generator is augmented with the verifier.
@@ A.5.2 Factors Effecting Performance on FlawedFictions
We investigate if length of a story has an effect on how accurately do different LLMs detect continuity errors in them by measuring correlation  [^11] We use Point-Biserial Correlation since CEEval-Full at an instance level is a discrete i.e. 0 or 1. between a story’s length (measured by counting number of words) and the CEEval-Full score on that story. We find negative correlation coefficients for all the models that we test and while the correlation values are low -0.1 to -0.2, for 13 out of 14 models the correlation observed is statistically significant (p-value << 0.05). Refer to the Table 7 for the exact values.
@@ A.9.1 Summarization
Summary generated with GPT-4o-mini Once upon a time, there was a poor woman living in a wretched hut, whose little boy ventured into the cold forest to gather wood. While wandering, he discovered a jagged white stone and, feeling pity for it, draped his jacket over it. Upon returning home without his jacket, his mother scolded him for his foolishness, insisting that charity begins at home. When the boy returned to the stone, he found it had turned over, revealing a chest filled with silver coins. Suspecting the money to be stolen, he tossed it into a nearby pond, keeping only a four-shilling piece that floated, believing it to be honest. The boy’s mother, furious at his actions, sent him away to seek his fortune. He wandered far and wide, but was deemed too small and weak for any work until he found a merchant who hired him to fetch wood and water. When the merchant prepared to journey to distant lands, the boy entrusted him with the four-shilling piece to buy something in return. The merchant, initially reluctant, ended up purchasing a cat from an old woman who wished to dispose of it. As the merchant sailed, a fierce storm struck, and upon reaching a strange land, he found a tavern filled with mice, where diners struggled to eat. Remembering the cat, he brought it forth, and the patrons were able to enjoy their meal in peace. They begged him to sell the cat, and he agreed for a hundred dollars. Afterward, the merchant encountered another storm and, upon reaching yet another land, sold the cat again for two hundred dollars, and then once more for three hundred dollars in a place plagued by larger rats. Throughout his journey, the merchant pondered how much of his earnings to give the boy, deciding initially to keep a portion for himself. However, when a terrible storm threatened the ship, he vowed to give the boy all the money. Upon reaching home safely, he honored his promise, presenting the boy with six hundred dollars and his daughter as well. The boy, once a mere scullion, became as wealthy as the merchant, living in splendor and happiness, and graciously took in his mother, declaring that charity does not solely begin at home.  Claude 3.5 Sonnet with Verfier’s Explanation of the Continuity Error There is a continuity error in the story regarding the amount of money the merchant earned from selling the cat. The merchant sold the cat three times for progressive amounts: 100,100,200, and 300,whichwouldtotal300,whichwouldtotal600. However, this creates a continuity error because once the cat was sold the first time, the merchant would no longer have possessed it to sell it two more times. Unless the story explicitly explained how the merchant regained possession of the cat between sales, this represents a logical impossibility in the narrative sequence.  Author’s Remark. The original story has the cat return to the ship every time a sale is done, a detail which is missing in the summary that creates the continuity error if cat was sold once, how the merchant managed to sell it again and again.